OpenStudy (lukecrayonz):

Which is the 25th term of the arithmetic sequence where a1 = 8 and a9 = 64 ? Which is the 25th term of the arithmetic sequence where a1 = 8 and a9 = 64 ? @Mathematics

OpenStudy (amistre64):

a9 = 8 + d(8) (64-8)/8 =d 8-1 = 7 an = 8 + 7(n-1)

OpenStudy (anonymous):

Try to break the numbers into smaller numbers, then you will see clearer the pattern.

OpenStudy (amistre64):

then: a25 = 8 + 7(24)

OpenStudy (lukecrayonz):

Yay <3 someone that remembers this :) Thank you so much.